In the expression (x tan α + y cot α) (x cot α +y tan α) - 4xy cot2 2α is?????
Pls send a detailed solution.....

Grade:11

1 Answers

Aditya Gupta
5 years ago
P= (x tan α + y cot α) (x cot α +y tan α) - 4xy cot2 2α = x^2+y^2+xy(tan^2α+cot^2α – 4cot2 2α)
now, we use the following identity: cotz-tanz=2cot2z
so, we have P= x^2+y^2+xy(tan^2α+cot^2α – (cotα – tanα)^2)=x^2+y^2+xy(2) = (x+y)^2